Find the Area of the largest round pizza that can be baked in an 14in by 18in cookie sheet. Can you please explain. Thank you
Answer by checkley77(12569)

The maximum circle would have a diameter of 12 in.
Thus the area of a 14 in pizza is:
A=pi*r^2
A=3.14(12/2)^2
A=3.14*6^2
A=3.14*36
A=113.04 in^2 ans.